OpenStudy (anonymous):

Estimate sqrt(99.8) using differentials (or equivalently, a linear approximation).

OpenStudy (anonymous):

Do you have a function we can use for a linear approximation?

OpenStudy (anonymous):

this is the entire question i'm not sure how to even start it

OpenStudy (anonymous):

Okay wait.

OpenStudy (anonymous):

Since it's a square root I assume f(x)= sqrt(x)

OpenStudy (anonymous):

Did you get that?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now take the derivative of sqrt(x) .

OpenStudy (anonymous):

1/2(x)^-1/2

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

The formula for a linearized function about a point x=a is : \[L(x) = f(a)+f'(a)(x-a)\]

OpenStudy (anonymous):

Since the question doesn't specify, we will assume the linearization occurs about x = 1.

OpenStudy (anonymous):

Do you follow?

OpenStudy (phi):

don't you think a= 100 is a better choice?

OpenStudy (anonymous):

Meh. Okay 100.

OpenStudy (anonymous):

I guess it's a better approximation.

OpenStudy (anonymous):

i guess a=100

OpenStudy (anonymous):

Now evaluate f(a) and f'(a)

OpenStudy (anonymous):

f(100)=10 f'(100)=1/20

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

now plug it into that equation I gave you :) .

OpenStudy (anonymous):

10+(1/20)(X-100)

OpenStudy (anonymous):

Now plug in 99.8 :) .

OpenStudy (anonymous):

9.99

OpenStudy (anonymous):

Which is VERY close to the actual value :) . That's your answer :) .

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

Welcome!